A University wishes to hire 4 faculty members for its newly established department. They have to select from a group of 8 Male and 7 Female applicants, all of whom are equally qualified to fill in the positions. If University selects 4 at random, what is the probability that:
a) All four will be Men
b) All four will be Women
c) Atleast one will be a Women
d) Atleast one will be a Men
1
Expert's answer
2021-09-08T08:24:04-0400

a)


"P(4\ Men)=\dfrac{\dbinom{8}{4}\dbinom{7}{0}}{\dbinom{7+8}{4}}""\dbinom{8}{4}=\dfrac{8!}{4!(8-4)!}=70"

"\dbinom{7}{0}=1"

"\dbinom{8+7}{4}=\dfrac{15!}{4!(15-4)!}=1365"


"P(4\ Men)=\dfrac{\dbinom{8}{4}\dbinom{7}{0}}{\dbinom{7+8}{4}}=\dfrac{70}{1365}=\dfrac{2}{39}"

b)


"P(4\ Women)=\dfrac{\dbinom{8}{0}\dbinom{7}{4}}{\dbinom{7+8}{4}}""\dbinom{8}{0}=1"

"\dbinom{7}{4}=\dfrac{7!}{4!(7-4)!}=35"

"\dbinom{8+7}{4}=\dfrac{15!}{4!(15-4)!}=1365"


"P(4\ Women)=\dfrac{\dbinom{8}{0}\dbinom{7}{4}}{\dbinom{7+8}{4}}=\dfrac{35}{1365}=\dfrac{1}{39}"

(c)


"P(at\ least\ 1\ Women)=1-P(0\ Women)"

"=1-P(4\ Men)=1-\dfrac{2}{39}=\dfrac{37}{39}"

d)


"P(at\ least\ 1\ Men)=1-P(0\ Men)"

"=1-P(4\ Women)=1-\dfrac{1}{39}=\dfrac{38}{39}"